# Korean Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

### v3

@@ Abstract
Korean Aspect-Based Sentiment Analysis via
@@ 1 Introduction
Figure 1: t-SNE visualization of last [CLS] embeddings extracted from KR3 test set by two different experimental BaselineXLM-R encoders. Our PL-cF encourages the encoder to produce discriminable representations of different sentiment polarities.
@@
In this study, we derive pseudo-labels for real Korean reviews using machine-translated English ABSA data, inspired by the past research Balahur and Turchi (2012); Hoshino et al. (2024). Moreover, we employ Dual filtering on the actual Korean corpus converted to implicit NLI task Hendrycks and Gimpel (2017); Sun et al. (2019); Feng et al. (2022), thereby constructing an effective framework coined as Korean ABSA using Pseudo-Classifier with Corpus Filtering (KPC-cF) for implicit-feature alignment. Through this process, we assess the impact of our constructed classifier on the practical alignment of actual reviews. We validate that the pseudo-classifier, generated through the sentence-pair approach, outperforms the single approach in translation task. Furthermore, using the model that predicts the translated dataset most effectively as a baseline, we generate pseudo-labels for actual data and conduct real-world testing of Korean ABSA. This involves subsequent fine-tuning the filtered corpus based on language-agnostic embedding similarity for review and aspect sentence pairs, along with setting a threshold for Maximum Softmax Probability (MSP) in pseudo-labels.
@@ 2.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur and Turchi Balahur and Turchi (2012) has shown that Machine Translation (MT) systems can obtain training data for languages other than English in general sentiment classification. However, existing research using translation data for alignment and alignment-free methods Li et al. (2020); Zhang et al. (2021) inadequately address the challenge of universal knowledge transfer for linguistic subgroups like Korean. Also, although it was a different domain at Zhou et al. Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature $\Phi$ between translated source data $\bm{D_{S}}$ and actual target language data $\bm{D_{T}}$. Therefore, we attempted the following two phases to assess the impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Fig. 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, similar to the findings of Hoshino et al. (2024), the optimal baseline model for Korean reasoning is identified from the pool of models trained and assessed utilizing the translation dataset. In Phase 2, we fine-tune the baseline model $\Psi_{pre}(\Phi(D_{S});\theta_{D_{S}})$, which was effective in training on $D_{S}$, by additionally incorporating pseudo-labeled actual $D_{T}$. Employing the tuned model $\Psi_{post}(\Phi(D_{T});\theta_{{D_{S}}\rightarrow D_{T}})$, we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, LaBSE and confidence score filtering are performed to enhance implicit features $\Phi(D_{T})$. Detailed description of our Language Adaptation for aligned ABSA task is provided in Appendix§B.
@@ 2.2 LaBSE based Filtering
Table 1: Number of instances and average scores (LaBSE, MSP in 4 classes) for KR3 fine-tuning set.
@@ 2.3 Confidence score Filtering
Storage/Docs. Precision Recall Micro-F1 4-way acc 3-way acc Binary
BaselinemBERT+PL 4.74MB 15.23K $D_{T}$ 89.71 77.60 83.21 77.96 80.15 84.01
BaselinemBERT+PL-cF 2.53MB 6.77K $D_{T}$ 87.45 76.28 81.46 78.49 80.69 84.70
BaselineXLM-R+PL 4.60MB 15.23K $D_{T}$ 90.58 79.26 84.54 83.36 85.71 89.88
BaselineXLM-R+PL-cF 2.15MB 6.08K $D_{T}$ 89.16 79.24 83.89 83.89 86.25 90.40
BaselinemBERT+TR+PL 6.28MB 30.45K $D_{S}\sqcup D_{T}$ 92.15 78.28 84.64 80.45 82.71 86.57
BaselinemBERT+TR+PL-cF 4.07MB 21.99K $D_{S}\sqcup D_{T}$ 90.15 81.20 85.42 80.88 83.15 86.96
BaselinemBERT+TR+PL 6.28MB 30.45K $D_{S}\rightarrow D_{T}$ 91.08 77.48 83.73 78.37 80.58 84.60
BaselinemBERT+TR+PL-cF 4.07MB 21.99K $D_{S}\rightarrow D_{T}$ 90.03 77.80 83.47 78.94 81.16 85.09
BaselineXLM-R+TR+PL 6.14MB 30.45K $D_{S}\sqcup D_{T}$ 91.88 83.66 87.57 83.43 85.78 89.34
KPC-cF 3.69MB 21.30K $D_{S}\sqcup D_{T}$ 92.72 83.30 87.75 83.52 85.90 89.56
BaselineXLM-R+TR+PL 6.14MB 30.45K $D_{S}\rightarrow D_{T}$ 91.50 84.23 87.69 84.21 86.57 90.35
KPC-cF 3.69MB 21.30K $D_{S}\rightarrow D_{T}$ 92.46 84.53 88.29 84.34 86.72 90.89
Table 2: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to total number (i.e., equivalent to Kor-SemEval) and capacity of samples from KR3 Train. The performances where filtered models outperform non-filtered ones under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Appendix§B.3.
@@ 4 Discussion
BaselineXLM-R+TR $D_{S}$ 4.53 1.18
BaselineXLM-R+ PL $D_{T}$0.8 2.59
BaselineXLM-R+TR+PL $D_{S}\rightarrow D_{T}$ 3.26 3.29
KPC-cF $D_{S}\rightarrow D_{T}$ 4.98 1.20
Table 3: Comparisons of intermediate layer features on general Korean STS. We report Spearman’s $\rho$ (%). Higher is better. KLUE-STS discussed in Appendix§E.
Looking at Figure 4, In BaselineXLM-R+TR, the Target None class exhibits relatively high softmax probabilities, while data with polarity show relatively low probabilities, indicating significant confidence in polarity and asymmetric learning between classes. The results of KPC-cF demonstrate that additional fine-tuning with filtered data significantly increases the maximum probability of the model for Target Polarity. Simultaneously, with a reduction of over 40% in data, it prevents biased class learning, forming similar distributions among Target classes. Additionally, for polarity data where the target is detected, the regression line shows that higher LaBSE scores correspond to increased confidence. Enhancing this characteristic of the model, our KPC-cF demonstrates an improved ability to detect general semantic relationships in Korean, as shown in Tab. 3. In this experiment, we generated intermediate embeddings for KorSTS ham-etal-2020-kornli sentence pairs in a zero-shot manner and measured their similarity, presenting the results in terms of Spearman’s correlation ($\rho$) with the Ground Truth similarities. On the other hand, by providing word level features as input to a low-complexity LR model, we further confirmed the efficacy of the basic characteristics of the filtered data through observed variations in training outcomes (see Tab. 4). Our framework suggests improving task stability and reliability by setting an optimal threshold to adjust biased classes or data with uncertain classification outcomes.
Model Joint TR+PL $\rightarrow$ $\Delta$ TR+PL-cF
Train Set Micro-F1 4-way acc
XLM-RBase I.T + LR +16.01 +6.29
mBERT I.T + LR +13.14 +5.19
TF-IDF I.T + LR +0.19 +0.19
Table 4: Performance variation of TR+PL vs. TR+PL-cF (generated by BaselineXLM-R) utilizing Logistic Regression as a shallow feature classifier for different Input Token extraction. We combined each fine-tuning set and trained all models.
@@ 4.1 Further Analysis
To analyze the types of anticipatory effects that enable language models to exhibit improved performance in filtering corpora, we conducted a qualitative analysis by examining 116 cases where KPC-cF outputs matched Ground Truth but not those of the misaligned model (i.e., BaselineXLM-R+TR+PL). Interestingly, we identified the following four patterns that complement the predictive behavior of TR+PL futher training: (1) False Detection Refining: TR+PL incorrectly predicted aspects not mentioned in the reviews with a specific polarity, but KPC-cF refined these to ‘None’. (2) Mismatch Refining: TR+PL incorrectly predicted aspects mentioned in the reviews with a completely different polarity, which KPC-cF refined to the correct polarity. (3) Multi Context Detecting: KPC-cF detected contexts within a review with multiple aspects, correctly assigning the ‘None’ overlooked by TR+PL to the correct polarity. (4) Single Context Detecting: KPC-cF detected contexts within a single-aspect review and correctly assigned the polarity that TR+PL had overlooked as ‘None’.
@@ 5 Conclusion
Aspect-Based Sentiment Analysis (ABSA) has been recognized as one of the most compelling subfields in text analytics and NLP. However, obtaining high quality labeled data has been one of the most important issues hindering the developments of ABSA. In this study, we addressed the language gap in ABSA by constructing a pseudo-classifier. This approach involved aligning implicit features through dual filtering and showcased its industrial and research potential via feature validation across various layers. Additionally, we presented Kor-SemEval, KR3 test (Gold) composed of Korean fine-grained set. We invite the community to extend Korean ABSA by providing new datasets, trained models, evaluation results, and metrics.
@@ Acknowledgements
We thank Chanseo Nam for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research internship. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
@@ B.1 Task description
B.2 Classification approach
Single sentence Classification
Sentence-pair Classification
Ensemble
B.3 Evaluation of Language Adaptation
@@ C.2 Hyperparameter
C.3 Metrics
Appendix D Preliminary Experiment
D.1 Experiment: Kor-SemEval
Results
@@ Appendix E Limitations and Future Work
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Tab. 9, 10). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None.
Model Task Type KLUE-STS-Mid.
Zero-shot Eval.
[…]
BaselineXLM-R+TR $D_{S}$9.0
BaselineXLM-R+ PL $D_{T}$ 3.24
BaselineXLM-R+TR+PL $D_{S}\rightarrow D_{T}$4.3
KPC-cF $D_{S}\rightarrow D_{T}$ 3.21
Table 11: Comparisons of intermediate layer features on KLUE-STS. We report Spearman’s $\rho$ (%). Higher is better. Data within the mid-range score were used.
Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. This limitation is evident not only in the qualitative analysis of specific data point errors but also in the KLUE-STS Park et al. (2021) results in Tab. 3. However, we have confirmed through various evaluations (Tab. 2, 4, 5, 10, 11 and Fig. 3) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations Haig (2003); Udomcharoenchaikit et al. (2022), we propose upper filtering that enables Target None predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs.
Thus, future research will delve into the association between embeddings from Bi-Encoder and Cross-Encoder models, and explore alternative methods for filtering or joint representation of translated NLI set. Simultaneously, the KR3 test set effectively capture real-world imbalances, deliberately selecting sentences with a broader vocabulary is necessary for enhanced diversity. Fully constructing the KR3 dataset through human annotation, it will serve as a benchmark for assessing LMs in Korean ABSA. Furthermore, Korean data from other domains should be acquired and evaluated.

### v4

# KPC-cF: Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

@@ Abstract
KPC-cF: Aspect-Based Sentiment Analysis via
@@ 1 Introduction
Figure 1: t-SNE visualization of last [CLS] embeddings extracted from KR3 test set by two experimental BaselineXLM-R encoders. Our PL-cF encourages the encoder to produce discriminable representations of different sentiment polarities, focusing on relevant aspects.
@@
In this study, we derive pseudo-labels for real Korean reviews using machine-translated English ABSA data, inspired by the past research Balahur and Turchi (2012); Hoshino et al. (2024). Moreover, we employ Dual filtering on the actual Korean corpus converted to implicit NLI task Hendrycks and Gimpel (2017); Sun et al. (2019); Feng et al. (2022), thereby constructing an effective framework coined as Korean ABSA using Pseudo-Classifier with corpus Filtering (KPC-cF) for implicit-feature alignment. Through this process, we assess the impact of our constructed framework on the practical alignment of actual reviews. We validate the pseudo-classifier, generated through the sentence-pair approach, and it outperforms the single approach in translation task. Furthermore, using the model that predicts the translated dataset most effectively as a baseline, we generate pseudo-labels for actual data and conduct real-world testing of Korean ABSA. This involves subsequent fine-tuning the filtered corpus based on language-agnostic embedding similarity for review and aspect sentence pairs, along with setting a threshold for Maximum Softmax Probability (MSP) in pseudo-labels.
@@ 2.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur and Turchi Balahur and Turchi (2012) has shown that machine translation systems can obtain training data for languages other than English in general sentiment classification. However, using translation data for alignment and alignment-free methods Li et al. (2020); Zhang et al. (2021) inadequately address the challenge of universal knowledge transfer for linguistic subgroups like Korean. Also, although it was a different domain at Zhou et al. Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature $\Phi$ between translated source data $\bm{D_{S}}$ and actual target language data $\bm{D_{T}}$. Therefore, we attempt the following two phases to assess impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Fig. 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, we investigate the data structure exhibiting performance aligned with $D_{S}$ by thoroughly learning the fine patterns of the labeled benchmark for each model. NLI structure is referenced based on our research domain and language settings, following the prior work Hoshino et al. (2024). In Phase 2, we fine-tune the baseline $\Psi_{pre}(\Phi(D_{S});\theta_{D_{S}})$, which was effective in training on $D_{S}$, by additionally incorporating pseudo-labeled $D_{T}$. Employing the tuned model $\Psi_{post}(\Phi(D_{T});\theta_{{D_{S}}\rightarrow D_{T}})$, we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, Dual filtering are performed to enhance implicit features $\Phi(D_{T})$. Detailed description of our task is provided in Appendix§B.
@@ 2.2 LaBSE based Filtering
Table 1: Number of instances and average scores (LaBSE, MSP in 4 classes) for KR3 fine-tuning set. We construct model-specific dataset, PL and PL-cF.
@@ 2.3 Confidence score Filtering
Storage/Tokens Precision Recall Micro-F1 4-way acc 3-way acc Binary
BaselinemBERT+PL 4.74MB 1.28M $D_{T}$ 89.71 77.60 83.21 77.96 80.15 84.01
BaselinemBERT+PL-cF 2.53MB 0.68M $D_{T}$ 87.45 76.28 81.46 78.49 80.69 84.70
BaselineXLM-R+PL 4.60MB 1.13M $D_{T}$ 90.58 79.26 84.54 83.36 85.71 89.88
BaselineXLM-R+PL-cF 2.15MB 0.52M $D_{T}$ 89.16 79.24 83.89 83.89 86.25 90.40
BaselinemBERT+TR+PL 6.28MB 1.73M $D_{S}\sqcup D_{T}$ 92.15 78.28 84.64 80.45 82.71 86.57
BaselinemBERT+TR+PL-cF 4.07MB 1.13M $D_{S}\sqcup D_{T}$ 90.15 81.20 85.42 80.88 83.15 86.96
BaselinemBERT+TR+PL 6.28MB 1.73M $D_{S}\rightarrow D_{T}$ 91.08 77.48 83.73 78.37 80.58 84.60
BaselinemBERT+TR+PL-cF 4.07MB 1.13M $D_{S}\rightarrow D_{T}$ 90.03 77.80 83.47 78.94 81.16 85.09
BaselineXLM-R+TR+PL 6.14MB 1.51M $D_{S}\sqcup D_{T}$ 91.88 83.66 87.57 83.43 85.78 89.34
KPC-cF 3.69MB 0.89M $D_{S}\sqcup D_{T}$ 92.72 83.30 87.75 83.52 85.90 89.56
BaselineXLM-R+TR+PL 6.14MB 1.51M $D_{S}\rightarrow D_{T}$ 91.50 84.23 87.69 84.21 86.57 90.35
KPC-cF 3.69MB 0.89M $D_{S}\rightarrow D_{T}$ 92.46 84.53 88.29 84.34 86.72 90.89
Table 2: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to the overall capacity of structured NLI data and total number of valid tokens from KR3 Train. The performances where filtered models outperform non-filtered ones under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Appendix§B.3.
@@ 4 Discussion
BaselineXLM-R+TR $D_{S}$ 10.26 2.55*
BaselineXLM-R+ PL $D_{T}$ 12.18 7.59
BaselineXLM-R+TR+PL $D_{S}\rightarrow D_{T}$ 9.21 6.35
KPC-cF $D_{S}\rightarrow D_{T}$ 16.26 20.50
Table 3: Comparisons of intermediate layer features as mean-pooled embeddings Choi et al. (2021) on general STS. We report Spearman’s $\rho$ (%). Higher is better. *Values are for p $\approx$ 0.1; others within p < 0.05. SetFit is selected based on the number of fine-tuning data. [^4] For details on SetFitKLUE-R, refer to Tunstall et al. (2022) and https://huggingface.co/mini1013.
Looking at Figure 4, In BaselineXLM-R+TR, the Target None class exhibits relatively high softmax probabilities, while data with polarity show relatively low probabilities, indicating significant confidence in polarity and asymmetric learning between classes. The results of KPC-cF demonstrate that additional fine-tuning with filtered data significantly increases the maximum probability of the model for Target Polarity. Simultaneously, with a reduction of over 40% in data, it prevents biased class learning, forming similar distributions among Target classes. Additionally, for polarity data where the target is detected, the regression line shows that higher LaBSE scores correspond to increased confidence. Meanwhile, our KPC-cF approach demonstrated more scalable performance than other training methods in comparisons of intermediate and shallow layer features. KPC-cF demonstrates an improved ability to detect general semantic relationships in Korean, as shown in Tab. 3. In this experiment, we generated zero-shot intermediate embeddings for KorSTS, KLUE-STS Ham et al. (2020); Park et al. (2021) sentence pairs, evaluating similarity with Spearman’s correlation ($\rho$) to Ground Truth. On the other hand, by providing word level features as input to a low-complexity LR model, we further confirmed the efficacy of basic characteristics of filtered data through variations in training outcomes (see Tab. 4). Our framework enhances stability and reliability by optimally thresholding to adjust biased or uncertain data.
Model Feature Span TR+PL $\rightarrow$ $\Delta$ TR+PL-cF
${D_{S\sqcup T}}$ ${x_{s\oplus a}}$ Micro-F1 4-way acc
XLM-RBase I.T + LR S.T +16.01 +6.29
mBERT I.T + LR S.T +13.14 +5.19
TF-IDF I.T + LR - +0.19 +0.19
Table 4: Performance variation of TR+PL vs. TR+PL-cF (generated by BaselineXLM-R) utilizing Logistic Regression as a shallow feature classifier for different Input Token (IDs). We combined each Data & NLI format w/w.o special tokens.
@@ 4.1 Further Analysis
To analyze the types of anticipatory effects that enable language models to exhibit improved performance in filtering corpora, we conducted a qualitative analysis by examining 116 cases where KPC-cF outputs matched Ground Truth but not those of the misaligned model (i.e., BaselineXLM-R+TR+PL). Interestingly, we identified the following four patterns that complement the predictive behavior of BaselineXLM-R+TR+PL: (1) False Detection Refining: misaligned model incorrectly predicted aspects not mentioned in the reviews with a specific polarity, but KPC-cF refined these to ‘None’. (2) Mismatch Refining: misaligned model incorrectly predicted aspects mentioned in the reviews with a completely different polarity, which KPC-cF refined to the correct polarity. (3) Multi Context Detecting: KPC-cF detected contexts within a review with multiple aspects, correctly assigning the ‘None’ overlooked by misaligned model to the correct polarity. (4) Single Context Detecting: KPC-cF detected contexts within a single-aspect review and correctly assigned the polarity that misaligned model had overlooked as ‘None’.
@@ 5 Conclusion
Aspect-Based Sentiment Analysis (ABSA) is a key subfield in text analytics, but the lack of high-quality labeled data hinders its industrial development. This study aims to address the language gap in ABSA. By aligning implicit features through dual filtering, we propose an method to align target languages solely through translation of existing benchmarks. This approach is expected to optimize tasks for both Korean and low-resource languages with limited pre-training on rich native data.
@@ Acknowledgements
We thank Chanseo Nam (Korea National University of Education) for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research internship. We also thank Guijin Son (Yonsei University), Kyeongpil Kang (Kangwon National University), Joosun Yum (KAIST), Youngjoo Jo (ETRI), and Chunghun Ha (Hongik University) for their valuable discussions. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
@@ B.1 Task description
B.2 Classification approach
Single sentence Classification
Sentence-pair Classification
Ensemble
B.3 Evaluation of Language Adaptation
@@ C.2 Hyperparameter
C.3 Metrics
Appendix D Preliminary Experiment
D.1 Experiment: Kor-SemEval
Results
@@ Appendix E Limitations and Future Work
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Tab. 10, 11). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None.
Model Task Type KLUE-STS-Mid.
Zero-shot Eval.
[…]
BaselineXLM-R+TR $D_{S}$9.0
BaselineXLM-R+ PL $D_{T}$ 3.24
BaselineXLM-R+TR+PL $D_{S}\rightarrow D_{T}$4.3
KPC-cF $D_{S}\rightarrow D_{T}$ 3.21
Table 9: Comparisons of intermediate layer features as [CLS] embeddings on KLUE-STS. We report Spearman’s $\rho$ (%). Contrasting positive and negative trends emerge. Data within the mid-range score were used.
Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. However, we have confirmed through various evaluations (Tab. 2, 3, 4, 5, 9, 11 and Fig. 3) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations Haig (2003); Udomcharoenchaikit et al. (2022), we propose upper filtering that enables Target None predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs. Thus, future research will delve into the association between embeddings from Bi-Encoder and Cross-Encoder models, and explore alternative methods for filtering or joint representation of translated NLI set. Simultaneously, the KR3 test set effectively capture real-world imbalances, deliberately selecting sentences with a broader vocabulary is necessary for enhanced diversity. Fully constructing the KR3 dataset through human annotation, it will serve as a benchmark for assessing LMs in Korean ABSA. Furthermore, Korean data from other domains should be acquired and evaluated.